# ms-Mamba: Multi-scale Mamba for Time-Series Forecasting

### v1

@@ Abstract
The problem of Time-series Forecasting is generally addressed by recurrent, Transformer-based and the recently proposed Mamba-based architectures. However, existing architectures generally process their input at a single temporal scale, which may be sub-optimal for many tasks where information changes over multiple time scales. In this paper, we introduce a novel architecture called Multi-scale Mamba (ms-Mamba) to address this gap. ms-Mamba incorporates multiple temporal scales by using multiple Mamba blocks with different sampling rates ($\Delta$s). Our experiments on many benchmarks demonstrate that ms-Mamba outperforms state-of-the-art approaches, including the recently proposed Transformer-based and Mamba-based models. Codes and models will be made available.
@@
XXX
@@ 1 Introduction
Time-series Forecasting (TSF) is the problem of predicting future values of a variable of interest, given its history. This fundamental problem used to be generally addressed using recurrent architectures [37, 11] and long-short term memory [19] or their variants [7, 14], see, e.g., [24, 27] for detailed surveys. Such models are inherently well-suited to the task due to their sequential information modeling abilities. The introduction of self-attention based architectures, a.k.a. Transformers [32], enabled attending to more informative patterns and correlations across time and provided significant improvements. However, Transformers’ quadratic computational complexity has been a limiting factor.
State Space Models (SSMs) [17, 31] are reported to provide a better balance between performance and computational complexity. Recently proposed architectures based on SSMs, namely, Mamba [16], offer the promise of on-par or better performance than Transformer-based alternatives while running significantly faster. This has led to the widespread use of Mamba or its derivatives across different domains [16, 43, 30, 41].
The use of a Mamba-based approach for TSF was recently explored by Wang et al.[34]. Wang et al.introduced an architecture, called S-Mamba, which used Mamba in both the forward and reverse directions for TSF. Wang et al.showed that this simple approach obtained state-of-the-art (SOTA) results on many TSF benchmarks, often providing significant gains over Transformer-based architectures.
Figure 1: (a) Mamba and its variations (S-Mamba) use a single time-scale while processing time-series data. (b) Our ms-Mamba processes its input at different time-scales to better capture signal at different scales.
Time-series data generally consist of signals of multiple temporal scales. To better capture and exploit the multi-scale nature of time-series data, the literature has introduced extensions over the conventional models; e.g., multi-scale recurrent architectures [8], multi-scale convolution [22] and multi-scale Transformers [47, 5].
In this paper, we introduce ms-Mamba, a Mamba-based architecture for multi-scale processing of time-series data. To be specific, by leveraging on the versatility of SSMs’ learnable sampling rate, we construct a block that consists of multiple SSMs with different independent or inter-related sampling rates. We show that our ms-Mamba performs better than Transformer-based and Mamba-based architectures on several datasets.
@@ 2.1.1 Transformer-based Models
Transformers, initially introduced by Vaswani et al.[32], have revolutionized tasks that involve sequence processing and generation, with their self-attention mechanism proving highly effective in capturing long-range dependencies. This architecture, originally designed for natural language processing, has since been adapted to time-series forecasting tasks [1], primarily due to its ability to model complex temporal relationships. Duong-Trung et al.[10] demonstrate the efficacy of Transformers in long-term multi-horizon forecasting, addressing the challenge of vanishing correlations over extended horizons.
@@
Recent studies have aimed to address the limitations of standard Transformers in time-series applications. Foumani et al.[12] propose enhanced positional encodings to improve the positional awareness of the Transformer architecture in multivariate time-series classification. Lim et al.[25] propose a Transformer architecture to make use of a complex mix of inputs. Wang et al.[33] present Graphformer, a model that replaces traditional convolutional layers with dilated convolutional layers, thereby improving the efficiency of capturing temporal dependencies across multiple variates in a graph-based framework.
@@
Despite their unprecedented successes in natural language processing tasks, Transformer models face several challenges when applied to other time-series domains. One key limitation is their content-based attention mechanism, which struggles to detect crucial temporal dependencies, particularly in cases where dependencies weaken over time or when strong seasonal patterns are present [38]. Additionally, Transformers suffer from the quadratic complexity of the attention mechanism, which increases computational costs and memory usage, for long input sequences [36].
@@
To address these issues, several studies have proposed modifications to the self-attention mechanism. For instance, Zhou et al.[48] introduce Informer that employs a sparsified self-attention operation to lower computational complexity and improve long-term forecasting efficiency. Similarly, Wu et al.[39] propose Autoformer, which relies on an auto-correlation-based self-attention to better capture temporal dependencies.
2.1.2 Linear Models
@@
Linear models are another popular approach in TSF due to their simplicity and efficiency [2]. Oreshkin et al.[29] propose a stacked MLP based architecture with residual links. Challu et al.[3] improve this architecture with multi-rate data sampling and hierarchical interpolation for effectively modeling extra long sequences. Zeng et al.[44] analyze Transformers for TSF and found that simple linear mappings can outperform Transformer models especially when the data has strong periodic patterns. Chen et al.[6] introduce another notable linear approach, TSMixer, which leverages an all-MLP architecture to efficiently incorporate cross-variate and auxiliary information. Zhang et al.[45] propose LightTS which is tailored towards efficiently handling very long input series in multivariate TSF. Wang et al.[35] propose time-series Multi-layer Perceptron (MLP), which improves forecasting performance by incorporating domain-specific knowledge into the MLP architecture.
@@
While linear models with MLPs are simpler architectures and faster compared to Transformer-based models, they face several limitations. These models generally struggle with non-linear dependencies and tend to underperform in scenarios involving highly volatile or non-stationary patterns [6]. Moreover, compared to Transformer-based models, linear architectures are less effective at capturing global dependencies. This limitation necessitates longer input sequences to achieve comparable forecasting performance, which can increase the computational cost [42].
@@ 2.2 Mamba Models
Mamba [16] is a recent sequence model based on State Space Models (SSMs) [17, 31]. Due to its promise of better efficiency-performance trade-off, Mamba has quickly attracted interest from researchers across different domains [30, 41]. Mamba’s ability to perform content-based reasoning in linear complexity to the sequence length with its hardware-aware algorithm, made it an attractive alternative to the Transformer models. Several works have explored its application to time-series forecasting tasks. Wang et al.[34] propose S-Mamba, which relies on a bidirectional Mamba layer to capture inter-variate dependencies and an MLP to extract temporal dependencies. Their model achieves SOTA performance while being faster than Transformer-based alternatives.
Figure 2: An overview of the proposed method. ms-Mamba processes the time-series data at different sampling rates to better capture the multi-scale nature of the input signal. This is achieved by processing and updating the embeddings with different sampling rates (SR).
@@ 3.1 Problem Formulation
Time-series forecasting is the problem of estimating future $F$ values $\mathbf{Y}_{t+1:t+F}=\{\mathbf{x}_{t+1},\mathbf{x}_{t+2},...,\mathbf{x}_{t+F}\}\in\mathbb{R}^{F\times D}$ of a multi-variate time-series data given its recent $L$ values $\mathbf{X}_{t-L:t}=\{\mathbf{x}_{t-L},\mathbf{x}_{t-L+1},...,\mathbf{x}_{t}\}\in\mathbb{R}^{L\times D}$ as input. The task is to find the mapping $f$:
$f:\mathbf{X}_{t-L:t}\rightarrow\mathbf{Y}_{t+1:t+F},$ (1)
@@ 3.2 State Space Models (SSMs)
SSMs [17, 31] are sequence models which use a latent state space representation for representing a mapping between a time-series input $x(t)$ and the output $y(t)$ (considering a single-variate setting to simplify notation):
@@
where $\hat{\textbf{A}},\hat{\textbf{B}},{\textbf{C}}$ are derived by the chosen sampling function (e.g., for zeroth-order hold sampling, $\hat{\textbf{A}}=\exp(\Delta\mathbf{A})$, $\hat{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A}-I)\cdot\Delta\mathbf{B}$ [16]).
SSMs have been recently extended to work more efficiently through the use of a convolutional approach [18] and more effectively through better initialization [13]. Moreover, by relating $\hat{\textbf{A}},\hat{\textbf{B}},{\textbf{C}}$ to the input, the Mamba model [16] provides comparable or better performance than its Transformer-based counterparts.
@@ 4.1 Embedding Layer
Following prior work [15, 26], we first transform the input time-series data through an embedding layer (Figure 2). Given an input sequence $\mathbf{X}\in\mathbb{R}^{L\times D}$ with $L$ time steps and $D$ variables, we apply a linear transformation along the temporal dimension:
@@ 4.2 Multi-scale Mamba Layer
where $\mathbf{E}^{l}$ is the output of the embedding layer at layer $l$. We explore three different strategies for obtaining $\Delta_{i}$:
@@
    where $\mathrm{Flatten}(\cdot)$ reshapes the input tensor $\mathbf{E}^{l}\in\mathbb{R}^{L\times D_{e}}$ into a vector of dimension $L\cdot D_{e}$, and $\mathrm{MLP}(\cdot)$ consists of two linear layers with a ReLU activation in between: $\mathrm{MLP}(\mathbf{x})=\mathbf{W}_{2}\max(0,\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})+\mathbf{b}_{2}$, mapping the flattened input to $n$ different sampling rates.
To improve the effectiveness of sequential processing, we employ our Multi-scale Mamba module in both directions as illustrated in Figure 2, following prior work (e.g., [15, 50]).
@@ 4.3 Normalization, Feed-Forward Network and Projection
After the last encoder block ($\mathbf{E}^{L}\in\mathbb{R}^{L\times D_{e}}$), a linear projection layer is applied to map the embedding dimension to prediction length to obtain the final prediction ($\hat{y}\in\mathbb{R}^{F\times D}$):
$\hat{y}=\mathrm{Linear(\mathbf{E}^{L})}.$ (11)
@@ 4.4 Training Objective
where $\hat{y}_{i,j}$ and $y_{i,j}$ are the predicted and ground truth values for the $i$-th time step and $j$-th variable, respectively; $F$ is the forecast horizon; and $D$ is the number of variables. The model parameters are optimized using the Adam optimizer [20] – see the Suppl. Mat. for more details about the training and experimental details.
@@ 5.1 Experimental Details
Datasets. To evaluate our proposed ms-Mamba, we conduct extensive experiments on thirteen real-world time-series forecasting benchmark datasets. The datasets are grouped into three categories for easier comparison. (1) Traffic-related datasets, which include Traffic [39] and PEMS [4]. The Traffic dataset consists of hourly road occupancy rates from the California Department of Transportation, consisting of data collected from 862 sensors on San Francisco Bay area freeways from January 2015 to December 2016. PEMS datasets are complex spatial-temporal datasets for California’s public traffic networks, includes four subsets (PEMS03, PEMS04, PEMS07, PEMS08), similar to SCINet. These traffic-related datasets has many periodic features. (2) ETT (Electricity Transformer Temperature) datasets [48], which contain load and oil temperature data from electricity transformers, collected between July 2016 and July 2018. This group includes four subsets: ETTm1, ETTm2, ETTh1, and ETTh2, which have fewer variables and show less regularity compared to traffic datasets. (3) Other datasets: Electricity [39], Exchange [39], Weather [39], and Solar-Energy [21]. The Electricity dataset includes the hourly electricity usage of 321 customers from 2012 to 2014. Solar-Energy dataset contains solar power generation data from 137 PV plants in Alabama in 2006, recorded at 10 minute resolution. The Weather dataset includes 21 meteorological indicators also recorded at 10 minute resolution from the Max Planck Biogeochemistry Institute’s Weather Station in 2020. Exchange dataset compiles daily exchange rates for eight countries from 1990 to 2016. The prior two datasets of this category contain many features most of which are periodic, the last two datasets have fewer primarily aperiodic features. See the Suppl. Mat for a summary of the datasets.
Models ms-Mamba ms-Mamba w/ Learnable $\Delta$ S-Mamba iTrans- former RLinear PatchTST Cross- former TiDE TimesNet DLinear FED- former Auto- former
Traffic 96 0.376 0.375 0.382 0.395 0.649 0.462 0.522 0.805 0.593 0.650 0.587 0.613
192 0.392 0.384 0.396 0.417 0.601 0.466 0.530 0.756 0.617 0.598 0.604 0.616
336 0.405 0.408 0.417 0.433 0.609 0.482 0.558 0.762 0.629 0.605 0.621 0.622
720 0.452 0.442 0.460 0.467 0.647 0.514 0.589 0.719 0.640 0.645 0.626 0.660
Avg 0.406 0.402 0.414 0.428 0.626 0.481 0.550 0.760 0.620 0.625 0.610 0.628
PEMS03 12 0.065 0.066 0.065 0.071 0.126 0.099 0.090 0.178 0.085 0.122 0.126 0.272
24 0.087 0.087 0.087 0.093 0.246 0.142 0.121 0.257 0.118 0.201 0.149 0.334
48 0.131 0.133 0.133 0.125 0.551 0.211 0.202 0.379 0.155 0.333 0.227 1.032
96 0.197 0.201 0.201 0.164 1.057 0.269 0.262 0.490 0.228 0.457 0.348 1.031
Avg 0.120 0.122 0.122 0.113 0.495 0.180 0.169 0.326 0.147 0.278 0.213 0.667
PEMS04 12 0.074 0.072 0.076 0.078 0.138 0.105 0.098 0.219 0.087 0.148 0.138 0.424
24 0.086 0.083 0.084 0.095 0.258 0.153 0.131 0.292 0.10 0.224 0.177 0.459
48 0.102 0.099 0.115 0.120 0.572 0.229 0.205 0.409 0.136 0.355 0.270 0.646
96 0.130 0.121 0.137 0.150 1.137 0.291 0.402 0.492 0.190 0.452 0.341 0.912
Avg 0.098 0.094 0.103 0.111 0.526 0.195 0.209 0.353 0.129 0.295 0.231 0.610
PEMS07 12 0.060 0.060 0.063 0.067 0.118 0.095 0.094 0.173 0.082 0.115 0.109 0.199
24 0.075 0.075 0.081 0.088 0.242 0.150 0.139 0.271 0.101 0.210 0.125 0.323
48 0.091 0.091 0.093 0.110 0.562 0.253 0.311 0.446 0.134 0.398 0.165 0.390
96 0.111 0.109 0.117 0.139 1.096 0.346 0.396 0.628 0.181 0.594 0.262 0.554
Avg 0.085 0.084 0.089 0.101 0.504 0.211 0.235 0.380 0.124 0.329 0.165 0.367
PEMS08 12 0.074 0.073 0.076 0.079 0.133 0.168 0.165 0.227 0.112 0.154 0.173 0.436
24 0.097 0.098 0.104 0.115 0.249 0.224 0.215 0.318 0.141 0.248 0.210 0.467
48 0.156 0.154 0.167 0.186 0.569 0.321 0.315 0.497 0.198 0.440 0.320 0.966
96 0.243 0.236 0.245 0.221 1.166 0.408 0.377 0.721 0.320 0.674 0.442 1.385
Avg 0.143 0.140 0.148 0.150 0.529 0.280 0.268 0.441 0.193 0.379 0.286 0.814
@@
Table 1: Experiment 1: Quantitative comparison between ms-Mamba and the existing methods on traffic-related datasets. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 12, 24, 48, 96 for PEMS and 96, 192, 336, 720 for Traffic. Top results are highlighted in bold while the second bests are underlined.
@@
Compared Methods. We compare our model with 10 state-of-the-art (SOTA) time-series forecasting models belonging to 4 different model families: (1) Mamba based models: S-Mamba [34]; (2) Transformer based models: iTransformer [26], PatchTST [28], Crossformer [46], FEDformer [49], Autoformer [39]; (3) Linear based models: TiDE [9], DLinear [44], RLinear [23]; and (4) Temporal Convolution based models: TimesNet [40]. The following provides a brief overview of these models:

### v2

@@ Abstract
keywords:
Time-series forecasting , Mamba , Multi-scale Mamba
journal: Neurocomputing
[…]
[ceng]organization=Dept. of Computer Eng. and ROMER Robotics Center, Middle East Technical University, state=Ankara, country=Turkey
[…]
[arch]organization=Dept. of Architecture and ROMER Robotics Center, Middle East Technical University, state=Ankara, country=Turkey
@@ 1 Introduction
Time-series Forecasting (TSF) is the problem of predicting future values of a variable of interest, given its history (Lim and Zohren, 2021; Miller et al., 2024; Nobrega and Oliveira, 2019; George et al., 2023). This fundamental problem used to be generally addressed using recurrent architectures (Williams and Zipser, 1989; Elman, 1990) and long-short term memory (Hochreiter and Schmidhuber, 1997) or their variants (Chung et al., 2014; Graves and Schmidhuber, 2005), see, e.g., (Lim and Zohren, 2021; Miller et al., 2024) for detailed surveys. Such models are inherently well-suited to the task due to their sequential information modeling abilities. The introduction of self-attention based architectures, a.k.a. Transformers (Vaswani et al., 2017), enabled attending to more informative patterns and correlations across time and provided significant improvements. However, Transformers’ quadratic computational complexity has been a limiting factor.
State Space Models (SSMs) (Gu et al., 2021a; Smith et al., 2022) are reported to provide a better balance between performance and computational complexity. Recently proposed architectures based on SSMs, namely, Mamba (Gu and Dao, 2023), offer the promise of on-par or better performance than Transformer-based alternatives while running significantly faster. This has led to the wide-spread use of Mamba or its derivatives across different domains (Gu and Dao, 2023; Yue and Li, 2024; Qu et al., 2024; Xu et al., 2024).
The use of a Mamba-based approach for TSF was recently explored by Wang et al. (2024c). Wang et al. introduced an architecture, called S-Mamba, which used Mamba in both the forward and reverse directions for TSF. Wang et al. showed that this simple approach obtained state-of-the-art (SOTA) results on many TSF benchmarks, often providing significant gains over Transformer-based architectures.
Figure 1: (a) Time-series data often contain information at multiple time-scales. Illustrated here is temperature from the Weather dataset Wu et al. (2021), averaged over different window lengths to highlight different scales. See B.3 for similar multi-scale visualizations from other datasets. (b) Mamba and its variations (S-Mamba) use a single time-scale while processing time-series data. (c) Our ms-Mamba processes its input at different time-scales to better capture signal at different scales.
Motivation: Why multi-scale TSF? Time-series data generally consist of signals of multiple temporal scales (see, e.g., Figure 1(a)). A temperature signal, for example, can have trends at scales of hours, days (day and night), weeks, months or years. To better capture and exploit the multi-scale nature of time-series data, the literature has introduced extensions over the conventional models; e.g., multi-scale recurrent architectures (Chung et al., 2017), multi-scale convolution (Li et al., 2024) and multi-scale Transformers (Zhang et al., 2024; Chen et al., 2024).
Our Solution. In this paper, we introduce ms-Mamba, a Mamba-based architecture for multi-scale processing of time-series data (Figure 1(c)), in contrast to standard Mamba-based models that process data at a single time-scale (Figure 1(b)). To be specific, by leveraging on the versatility of SSMs’ learnable sampling rate, we construct a block that consists of multiple SSMs with different independent or inter-related sampling rates. We show that our ms-Mamba performs better than Transformer-based and Mamba-based architectures on several datasets.
@@ 2.1 Time-series Forecasting
Transformer-based Models. Transformers, initially introduced by Vaswani et al. (2017), have revolutionized tasks that involve sequence processing and generation, with their self-attention mechanism proving highly effective in capturing long-range dependencies. This architecture, originally designed for natural language processing, has since been adapted to time-series forecasting tasks (Ahmed et al., 2023), primarily due to its ability to model complex temporal relationships. Duong-Trung et al. (2023) demonstrate the efficacy of Transformers in long-term multi-horizon forecasting, addressing the challenge of vanishing correlations over extended horizons.
@@
Recent studies have aimed to address the limitations of standard Transformers in time-series applications. Foumani et al. (2024) propose enhanced positional encodings to improve the positional awareness of the Transformer architecture in multivariate time-series classification. Lim et al. (2021) propose a Transformer architecture to make use of a complex mix of inputs. Wang et al. (2024a) present Graphformer, a model that replaces traditional convolutional layers with dilated convolutional layers, thereby improving the efficiency of capturing temporal dependencies across multiple variates in a graph-based framework.
@@
Despite their unprecedented successes in natural language processing tasks, Transformer models face several challenges when applied to other time-series domains. One key limitation is their content-based attention mechanism, which struggles to detect crucial temporal dependencies, particularly in cases where dependencies weaken over time or when strong seasonal patterns are present (Woo et al., 2022). Additionally, Transformers suffer from the quadratic complexity of the attention mechanism, which increases computational costs and memory usage, for long input sequences (Wen et al., 2022).
@@
To address these issues, several studies have proposed modifications to the self-attention mechanism. For instance, Zhou et al. (2021) introduce Informer that employs a sparsified self-attention operation to lower computational complexity and improve long-term forecasting efficiency. Similarly, Wu et al. (2021) propose Autoformer, which relies on an auto-correlation-based self-attention to better capture temporal dependencies.
@@
Linear Models. Linear models are another popular approach in TSF due to their simplicity and efficiency (Benidis et al., 2022). Oreshkin et al. (2019) propose a stacked MLP based architecture with residual links. Challu et al. (2023) improve this architecture with multi-rate data sampling and hierarchical interpolation for effectively modeling extra long sequences. Zeng et al. (2023) analyze Transformers for TSF and found that simple linear mappings can outperform Transformer models especially when the data has strong periodic patterns. Chen et al. (2023) introduce another notable linear approach, TSMixer, which leverages an all-MLP architecture to efficiently incorporate cross-variate and auxiliary information. Zhang et al. (2022) propose LightTS which is tailored towards efficiently handling very long input series in multivariate TSF. Wang et al. (2024b) propose time-series Multi-layer Perceptron (MLP), which improves forecasting performance by incorporating domain-specific knowledge into the MLP architecture.
@@
While linear models with MLPs are simpler architectures and faster compared to Transformer-based models, they face several limitations. These models generally struggle with non-linear dependencies and tend to underperform in scenarios involving highly volatile or non-stationary patterns (Chen et al., 2023). Moreover, compared to Transformer-based models, linear architectures are less effective at capturing global dependencies. This limitation necessitates longer input sequences to achieve comparable forecasting performance, which can increase the computational cost (Yi et al., 2024).
Multi-scale Models. The literature has witnessed many multi-scale time-series models using different architectures. For instance, exploiting the Transformer architecture, Pyraformer (Liu et al., 2022b) constructs a multi-resolution pyramidal graph to capture long-range dependencies, while Crossformer (Zhang and Yan, 2023) utilizes Dimension-Segment-Wise embeddings to explicitly model information at varying granularities. Beyond pure attention, TimesNet (Wu et al., ) transforms 1D series into 2D to capture intra- and inter-period variations, and convolutional approaches like MICN (Wang et al., 2023) and SCINet (Liu et al., 2022a) leverage multi-branch convolutions or recursive downsampling to fuse local and global contexts. While these methods achieve multi-scale processing through explicit structural hierarchies or downsampling operations, our approach extends this capability to Mamba by leveraging the inherent discretization properties of the sampling rate $\Delta$, enabling multi-resolution feature extraction without architectural downsampling.
@@ 2.2 Mamba Models
Mamba Gu and Dao (2023) is a recent sequence model based on State Space Models (SSMs) Gu et al. (2021a); Smith et al. (2022). Due to its promise of better efficiency-performance trade-off, Mamba has quickly attracted interest from researchers across different domains Qu et al. (2024); Xu et al. (2024). Mamba’s ability to perform content-based reasoning in linear complexity to the sequence length with its hardware-aware algorithm, made it an attractive alternative to the Transformer models. Several works have explored its application to time-series forecasting tasks. Wang et al. (2024c) propose S-Mamba, which relies on a bidirectional Mamba layer to capture inter-variate dependencies and an MLP to extract temporal dependencies. Their model achieves SOTA performance while being faster than Transformer-based alternatives.
Figure 2: An overview of the proposed method. ms-Mamba processes the time-series data at different sampling rates to better capture the multi-scale nature of the input signal. This is achieved by processing and updating the embeddings with different sampling rates (SR). In the input visualization, each square represents a data point for a single variate at a specific time step, with the axes representing the Variates and Time dimensions, respectively.
@@ 3.1 Problem Formulation
Time-series forecasting is the problem of estimating future $T$ values $\mathbf{Y}_{t+1:t+T}=\{\mathbf{x}_{t+1},\mathbf{x}_{t+2},...,\mathbf{x}_{t+T}\}\in\mathbb{R}^{F\times D}$ of a multi-variate time-series data given its recent $L$ values $\mathbf{X}_{t-L:t}=\{\mathbf{x}_{t-L},\mathbf{x}_{t-L+1},...,\mathbf{x}_{t}\}\in\mathbb{R}^{L\times D}$ as input. The task is to find the mapping $f$:
$f:\mathbf{X}_{t-L:t}\rightarrow\mathbf{Y}_{t+1:t+T},$ (1)
@@ 3.2 State Space Models (SSMs)
SSMs Gu et al. (2021a); Smith et al. (2022) are sequence models which use a latent state space representation for representing a mapping between a time-series input $x(t)$ and the output $y(t)$ (considering a single-variate setting to simplify notation):
@@
where $\hat{\textbf{A}},\hat{\textbf{B}},{\textbf{C}}$ are derived by the chosen sampling function (e.g., for zeroth-order hold sampling, $\hat{\textbf{A}}=\exp(\Delta\mathbf{A})$, $\hat{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A}-I)\cdot\Delta\mathbf{B}$ Gu and Dao (2023)).
SSMs have been recently extended to work more efficiently through the use of a convolutional approach Gu et al. (2021b) and more effectively through better initialization Fu et al. (2022). Moreover, by relating $\hat{\textbf{A}},\hat{\textbf{B}},{\textbf{C}}$ to the input, the Mamba model Gu and Dao (2023) provides comparable or better performance than its Transformer-based counterparts.
@@ 4 Methodology: ms-Mamba
Influence of Sampling Rate on Temporal Resolution. By employing multiple Mamba blocks, each parameterized with a different sampling rate ($\Delta$), our model is designed to intrinsically capture features at multiple, distinct temporal resolutions (or time scales): Overall, large $\Delta$ induces shorter memory and lower temporal resolution whereas small $\Delta$ facilitates long-memory, high-resolution behavior. See Section A for a more in-depth argument.
@@ 4.1 Embedding Layer
Following prior work Grazzi et al. (2024); Liu et al. (2023), we first transform the input time-series data through an embedding layer (Figure 2). Given an input sequence $\mathbf{X}\in\mathbb{R}^{L\times D}$ with $L$ time steps and $D$ variables, we apply a linear transformation along the temporal dimension:
@@ 4.2 Multi-scale Mamba Layer
where $\mathbf{E}^{l}$ is the output of the embedding layer at layer $l$.
Note that each mamba block contains an internal normalization layer, so our averaging operation fuses the already-normalized outputs from each scale. Furthermore, keep in mind that the sampling rate $\Delta$ is an internal SSM parameter and not an input downsampling operation. All parallel Mamba blocks receive the same input $\mathbf{E}^{l}$ and produce identically-sized outputs, allowing for direct fusion without any temporal alignment.
We explore three different strategies for obtaining $\Delta_{i}$:
@@
    where $\mathrm{Flatten}(\cdot)$ reshapes the input tensor $\mathbf{E}^{l}\in\mathbb{R}^{L\times D_{e}}$ into a vector of dimension $L\cdot D_{e}$, and $\mathrm{MLP}(\cdot)$ consists of two linear layers with a ReLU activation in between: $\mathrm{MLP}(\mathbf{x})=\mathbf{W}_{2}\max(0,\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})+\mathbf{b}_{2}$, mapping the flattened input ($\mathbf{x}$, representing $\mathrm{Flatten}(\mathbf{E}^{l})$ in Equation 9) to $n$ different sampling rates.
To improve the effectiveness of sequential processing, we employ our Multi-scale Mamba module in both directions as illustrated in Figure 2, following prior work (e.g., Grazzi et al. (2024); Zhu et al. (2024)).
By using different sampling rates $\Delta_{i}$, each Mamba block generates different discrete-time SSM parameters, allowing each block to specialize in capturing temporal dependencies at a distinct scale. Section A for a more in-depth argument.
@@ 4.3 Normalization, Feed-Forward Network and Projection
After the last ($N^{th}$) encoder block ($\mathbf{E}^{N}\in\mathbb{R}^{L\times D_{e}}$), a linear projection layer is applied to map the embedding dimension to prediction length to obtain the final prediction ($\hat{y}\in\mathbb{R}^{F\times D}$):
$\hat{y}=\mathrm{Linear(\mathbf{E}^{N})}.$ (11)
@@ 4.4 Training Objective
where $\hat{y}_{i,j}$ and $y_{i,j}$ are the predicted and ground truth values for the $i$-th time step and $j$-th variable, respectively; $F$ is the forecast horizon; and $D$ is the number of variables. The model parameters are optimized using the Adam optimizer Kingma (2014) – see the Suppl. Mat. for more details about the training and experimental details.
@@ 5 Experiments
Unless otherwise stated, ms-Mamba refers to ms-Mamba with learnable temporal scales.
@@ 5.1 Experimental Details
Datasets. To evaluate our proposed ms-Mamba, we conduct extensive experiments on thirteen real-world time-series forecasting benchmark datasets. The datasets are grouped into three categories for easier comparison. (1) Traffic-related datasets, which include Traffic Wu et al. (2021) and PEMS Chen et al. (2001). The Traffic dataset consists of hourly road occupancy rates from the California Department of Transportation, consisting of data collected from 862 sensors on San Francisco Bay area freeways from January 2015 to December 2016. PEMS datasets are complex spatial-temporal datasets for California’s public traffic networks, includes four subsets (PEMS03, PEMS04, PEMS07, PEMS08), similar to SCINet. These traffic-related datasets have many periodic features. (2) ETT (Electricity Transformer Temperature) datasets Zhou et al. (2021), which contain load and oil temperature data from electricity transformers, collected between July 2016 and July 2018. This group includes four subsets: ETTm1, ETTm2, ETTh1, and ETTh2, which have fewer variables and show less regularity compared to traffic datasets. (3) Other datasets: Electricity Wu et al. (2021), Exchange Wu et al. (2021), Weather Wu et al. (2021), and Solar-Energy Lai et al. (2018). The Electricity dataset includes the hourly electricity usage of 321 customers from 2012 to 2014. Solar-Energy dataset contains solar power generation data from 137 PV plants in Alabama in 2006, recorded at 10 minute resolution. The Weather dataset includes 21 meteorological indicators also recorded at 10 minute resolution from the Max Planck Biogeochemistry Institute’s Weather Station in 2020. Exchange dataset compiles daily exchange rates for eight countries from 1990 to 2016. The prior two datasets of this category contain many features most of which are periodic, the last two datasets have fewer primarily aperiodic features.
See the Suppl. Mat for more details on the datasets.
@@
Table 1: Experiment 1: Quantitative results on traffic-related datasets. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 12, 24, 48, 96 for PEMS and 96, 192, 336, 720 for Traffic. Top results are highlighted in bold while the second bests are underlined.
Dataset $T$ ms-Mamba (Ours) S-Mamba iTransformer RLinear PatchTST Crossformer TiDE TimesNet DLinear FEDformer Autoformer
Traffic 96 0.375 0.382 0.395 0.649 0.462 0.522 0.805 0.593 0.650 0.587 0.613
192 0.384 0.396 0.417 0.601 0.466 0.530 0.756 0.617 0.598 0.604 0.616
336 0.408 0.417 0.433 0.609 0.482 0.558 0.762 0.629 0.605 0.621 0.622
720 0.442 0.460 0.467 0.647 0.514 0.589 0.719 0.640 0.645 0.626 0.660
[.4pt/1pt] Avg 0.402 0.414 0.428 0.626 0.481 0.550 0.760 0.620 0.625 0.610 0.628
PEMS03 12 0.066 0.065 0.071 0.126 0.099 0.090 0.178 0.085 0.122 0.126 0.272
24 0.087 0.087 0.093 0.246 0.142 0.121 0.257 0.118 0.201 0.149 0.334
48 0.133 0.133 0.125 0.551 0.211 0.202 0.379 0.155 0.333 0.227 1.032
96 0.201 0.201 0.164 1.057 0.269 0.262 0.490 0.228 0.457 0.348 1.031
[.4pt/1pt] Avg 0.122 0.122 0.113 0.495 0.180 0.169 0.326 0.147 0.278 0.213 0.667
PEMS04 12 0.072 0.076 0.078 0.138 0.105 0.098 0.219 0.087 0.148 0.138 0.424
24 0.083 0.084 0.095 0.258 0.153 0.131 0.292 0.10 0.224 0.177 0.459
48 0.099 0.115 0.120 0.572 0.229 0.205 0.409 0.136 0.355 0.270 0.646
96 0.121 0.137 0.150 1.137 0.291 0.402 0.492 0.190 0.452 0.341 0.912
[.4pt/1pt] Avg 0.094 0.103 0.111 0.526 0.195 0.209 0.353 0.129 0.295 0.231 0.610
PEMS07 12 0.060 0.063 0.067 0.118 0.095 0.094 0.173 0.082 0.115 0.109 0.199
24 0.075 0.081 0.088 0.242 0.150 0.139 0.271 0.101 0.210 0.125 0.323
48 0.091 0.093 0.110 0.562 0.253 0.311 0.446 0.134 0.398 0.165 0.390
96 0.109 0.117 0.139 1.096 0.346 0.396 0.628 0.181 0.594 0.262 0.554
[.4pt/1pt] Avg 0.084 0.089 0.101 0.504 0.211 0.235 0.380 0.124 0.329 0.165 0.367
PEMS08 12 0.073 0.076 0.079 0.133 0.168 0.165 0.227 0.112 0.154 0.173 0.436
24 0.098 0.104 0.115 0.249 0.224 0.215 0.318 0.141 0.248 0.210 0.467
48 0.154 0.167 0.186 0.569 0.321 0.315 0.497 0.198 0.440 0.320 0.966
96 0.236 0.245 0.221 1.166 0.408 0.377 0.721 0.320 0.674 0.442 1.385
[.4pt/1pt] Avg 0.140 0.148 0.150 0.529 0.280 0.268 0.441 0.193 0.379 0.286 0.814
@@
Compared Methods. We compare our model with 10 state-of-the-art (SOTA) time-series forecasting models belonging to 4 different model families: (1) Mamba based models: S-Mamba Wang et al. (2024c); (2) Transformer based models: iTransformer Liu et al. (2023), PatchTST Nie et al. (2022), Crossformer Zhang and Yan (2023), FEDformer Zhou et al. (2022), Autoformer Wu et al. (2021); (3) Linear based models: TiDE Das et al. (2023), DLinear Zeng et al. (2023), RLinear Li et al. (2023); and (4) Temporal Convolution based models: TimesNet Wu et al. . The following provides a brief overview of these models:
  • 1.